# Towards Solving the Gilbert-Pollak Conjecture via Large Language Models

### v1

@@ 1 Introduction
Recently, Large Language Models (LLMs) have demonstrated remarkable results in solving contest-level mathematics (Achim et al., ; Chen et al., 2025; Chervonyi et al., 2025). A natural next step is to explore whether these capabilities can be extended to tackling unsolved mathematical conjectures, such as the Gilbert-Pollak Conjecture. However, naïvely prompting LLMs to address such tasks often yields incorrect proofs with subtle logical gaps and fundamental misunderstandings of key concepts. Unlike canonical definitions commonly found in textbooks, research-level concepts appear only sparsely in training data. Compared to well-structured competition problems, open research problems exhibit a vastly larger and less constrained search space. As a result, models struggle to identify valid proof strategies and become substantially more prone to hallucination.
@@
Reproducible codes and the final proof can be found in https://github.com/keyisi2006/Steiner-Ratio.
2 Related Works
Solving Math Problems with LLMs. Recent advancements in LLMs have demonstrated remarkable potential in mathematical reasoning, extending from standard benchmarks to competition-level challenges. In specialized domains, Trinh et al. (2024) introduced AlphaGeometry, a neuro-symbolic system capable of solving IMO geometry problems without human demonstrations. To tackle broader formal reasoning, HyperTree Proof Search (Lample et al., 2022) pioneered the integration of Transformer-based policies with Monte-Carlo Tree Search (MCTS). Building on this direction, recent works leverage Reinforcement Learning (RL) within formal environments: AlphaProof (Google DeepMind, 2024) achieves silver-medal standards in Lean, while DeepSeek-Prover (Xin et al., 2024a, b) employs RL to establish state-of-the-art performance in open-ended theorem proving. To enhance reasoning reliability, InternLM-Math (Ying et al., 2024) leverages process supervision to verify intermediate steps. On the infrastructure front, Yang et al. (2023) developed LeanDojo, an open environment enabling large-scale training of these systems. Most recently, Goedel-Prover (Lin et al., 2025a, b) was introduced as a frontier model, achieving top performance among open-weight models on the MiniF2F benchmark.
AI for Scientific Discovery. Artificial Intelligence has catalyzed paradigm shifts across various scientific disciplines. In biology, AlphaFold (Jumper et al., 2021, 2020; Krokidis et al., 2025) revolutionized structural biology by predicting protein structures with atomic accuracy. In materials science, Merchant et al. (2023) leveraged graph neural networks to discover 2.2 million new crystals (GNoME), dramatically expanding the scale of stable materials known to humanity. Similarly, GraphCast (Lam et al., 2023) demonstrated that AI can model complex physical dynamics to outperform traditional weather forecasting. Moving towards algorithmic discovery, AlphaTensor (Fawzi et al., 2022) utilized reinforcement learning to discover matrix multiplication algorithms surpassing human-designed heuristics. More recently, the focus has shifted towards fully automated research frameworks; for instance, Lu et al. (2024); Yamada et al. (2025) proposed ”The AI Scientist,” where LLM-based agents independently generate ideas, execute experiments, and write manuscripts. Following this trajectory, we extend the exploratory capabilities of AI to the domain of Theoretical Mathematics, specifically targeting combinatorial optimization problems.
@@ 3.2 Mathematical Framework for Proving Steiner Ratio Lower Bound
In this section, we briefly describe the mathematical framework for proving lower bound for Steiner ratio. Many of the mathematical details can be found in Appendix A.1.1.
We first give a very brief and informal description. The framework for establishing the Steiner ratio lower bound is mathematical induction. Assuming that for any set of points with cardinality at most n1n-1, the Steiner ratio is lower-bounded by $\rho$. To extend this to a set $V$ of $n$ points, the core strategy involves a reduction argument: identifying a specific subset $V^{\prime}\subset V$ of size nrn-r (where $r\geq 1$) such that the marginal gain in the Steiner tree length satisfies $\Delta L_{S}/\Delta L_{m}\geq\rho$. Here, $\Delta L_{S}=L_{S}(V)-L_{S}(V^{\prime})$ and $\Delta L_{m}=L_{m}(V)-L_{m}(V^{\prime})$ denote the reduction in the lengths of the Steiner Minimal Tree and the Minimum Spanning Tree, respectively. Since $|V^{\prime}|\leq n-1$, the inductive hypothesis guarantees $L_{S}(V^{\prime})/L_{m}(V^{\prime})\geq\rho$. Combining this with the marginal condition $\Delta L_{S}\geq\rho\Delta L_{m}$ completes the inductive step, proving $L_{S}(V)/L_{m}(V)\geq\rho$.
@@
Figure 4: Illustration of Pruning Process. We try to prune $V^{*}=\{B\}$. Black dashed lines are $S^{-}$. Black solid lines are $S^{+}$. Green line is $t^{*}$.
In this pruning process, the decrease in the Steiner tree length is given by $\Delta L_{S}=L_{S^{-}}-L_{S^{+}}$, while the increase in the spanning tree length is given by $\Delta L_{m}=L_{t^{*}}$. Here, $L_{X}$ denotes the total edge length of graph $X$. The requirement for the induction to hold is that $\Delta L_{S}/\Delta L_{m}\geq\rho$, which is equivalent to $\rho\cdot L_{t^{*}}+L_{S^{+}}-L_{S^{-}}\leq 0$. If this condition holds, then combining it with the inductive hypothesis $L_{S}(V^{\prime})/L_{m}(V^{\prime})\geq\rho$ allows us to derive $L_{S}(V)/L_{m}(V)\geq\rho$, thereby completing the induction step. Therefore, we can reformulate the verification of a candidate lower bound $\rho$ for the Steiner ratio as a minimax condition as Theorem 4. In the following sections, we introduce an LLM-based mathematical framework designed to address this minimax problem.
@@
Let $\mathcal{W}$ denote the space of all edge-length vectors associated with the tree topology. For any choice of splitting $\tau=(V^{*},S^{-},S^{+},t^{*})$ and geometric configuration ${\bm{w}}\in\mathcal{W}$, we define the splitting function $F_{\tau}({\bm{w}},\rho)$ as: $F_{\tau}({\bm{w}},\rho)=\rho\cdot L_{t^{*}}+L_{S^{+}}-L_{S^{-}}.\vskip-5.0pt$ Let $\mathcal{F}$ be a family of functions corresponding to admissible splits $\tau$, which we denote as spliting function. The lower bound $\rho_{\text{Steiner}}\geq\rho$ is established if: $\textstyle\max_{{\bm{w}}\in\mathcal{W}}\min_{F\in\mathcal{F}}F({\bm{w}},\rho)\leq 0.$ (1) We say that a value $\rho$ is feasible if the condition equation 1 holds.
@@ 4 LLM Math System for the Conjecture
Finally, we develop a bottleneck region-based reflection mechanism. Leveraging the reward signal and the verification process, we introduce an algorithm that automatically identifies key bottleneck regions limiting further improvements of the lower bound and feeds this information back to the LLM in natural language, guiding subsequent generations to overcome these bottlenecks. Experiments show that this signal significantly improves the effectiveness of LLM, thereby addressing Challenge (3).
@@ 4.1 Reward Model
The general objective of the reward in LLM is to evaluate the generated output. In our context, the reward is to determine the tightest lower bound for the Steiner ratio from the “generated proof”. Theorem 4 establishes that the statement “the Steiner ratio is lower bounded by $\rho$” is equivalent to verifying that $\max_{{\bm{w}}\in\mathcal{W}}\min_{F\in\mathcal{F}}F({\bm{w}},\rho)\leq 0$. However, directly evaluating this continuous-domain minimax problem is computationally intractable.
@@
A function $f:\mathcal{W}\to\mathbb{R}\cup\{+\infty\}$ is called a verification function if it satisfies the following two conditions: 1. Shape Constraint: The restriction of $f$ to any line parallel to a coordinate axis is unimodal (first non-increasing and then non-decreasing). 2. Bounding Constraint: There exists a splitting function $g\in\mathcal{F}$ such that $g({\bm{w}})\leq f({\bm{w}})$, $\forall{\bm{w}}\in\mathcal{W}$. We denote the set of all verification functions by $\mathcal{F}_{\text{ver}}$.
@@
Let $H\subset\mathcal{W}$ be a bounded axis-aligned hyperrectangle and let $f$ be any function in $\mathcal{F}_{\text{ver}}$. A key property is that the global maximum of $f$ over $H$ is attained at one of its vertices. Consequently, if $\textstyle\max_{{\bm{w}}\in\mathcal{V}(H)}f({\bm{w}})\leq 0,$ (2) then $f({\bm{w}})\leq 0$ for all ${\bm{w}}\in H$, which implies that the underlying splitting function $g({\bm{w}})\leq 0$.
Leveraging these theoretical properties of the verification function, we construct a reward model to compute the lower bound of the Steiner ratio. We present the core algorithm below, deferring the full pseudocode and detailed explanations to Appendix D.1.
@@
(1) Region Partitioning. We recursively partition the infinite parameter space into hyperrectangle regions. (2)Vertex Verification. For bounded regions, we select a verification function $f$ and invoke the Vertex Maximum Property. By verifying $f({\bm{w}})\leq 0$ at all vertices, we mathematically guarantee that the inequality holds for every point $\bm{w}$} within the region. (3) Monotonicity Check for Unbounded Regions. For unbounded regions, we first verify that $f$ is non-increasing along those unbounded dimensions, ensuring the maximum value occurs at the finite lower boundary (a “slice” of the region). We then apply vertex verification to this bounded slice. Details of the monotonicity check are provided in Appendix B.4. (4) Refinement via Subdivision. Unverified regions are subdivided along one dimension into two halves, and recursively attempt verification.
@@ 4.2 LLM Agent
We demonstrate that finding a verification function reduces to identifying two types of simpler geometric lemmas: (1) the Trapped Regular Point Lemma and (2) the Valid 4-Point Steiner Tree Lemma. (This equivalence is illustrated in Figure 8 and formally proven in Theorem 18, Appendix B.2). Within our framework, the LLM generates candidate lemmas in the form of structured code snippets. This approach significantly simplifies the generation process and facilitates rigorous verification. In the subsequent sections, we detail how we leverage the LLM to discover these lemmas.
@@ 4.2.1 Finding Trapped Regular Point Lemma
In this section, we detail the methodology for discovering Trapped Regular Point Lemmas using LLMs. We define two distinct categories of rules: Type A rules encode necessary properties that any Steiner Minimal Tree (SMT) must satisfy, while Type B rules specify sufficient conditions that guarantee the existence of an implicit regular point. Given the vast combinatorial space of rule interactions—which renders manual exploration intractable (see Appendix F for details)—we employ an LLM integrated with local tools (e.g., Mathematica). By iteratively validating hypotheses, the LLM discovers effective rule combinations and synthesizes conditions that imply regularity. These conditions are subsequently simplified by the model to derive novel geometric lemmas.
@@
These predicates correspond to algebraic inequalities on $\bm{w}$} (e.g., $C^{B}_{0,2,1}\iff(a_{2}\leq 1)\land(\sqrt{3}/2\cdot(1+a_{2})\leq 1)$). Full derivations are in Appendix B.1.
@@
LLM-Guided Lemma Generation. This module operationalizes Theorem 9 through the synthesis of valid constraints via an LLM agent. Here, we delineate the process of guiding the LLM to identify trapped regular point lemmas. Comprehensive details are available in Appendix F.
@@ 4.2.2 Finding Valid 4-Point Steiner Tree Lemma
We adopt the Valid 4-Point Steiner Tree Theorem from Du et al. (1987) to determine the existence of the (AB)(AB)-$(CD)$ topology. The theorem’s details are provided in Appendix A.2 as Theorem 14. The LLM agent’s task is to translate the geometric conditions of the theorem into algebraic constraints on the parameter space $\bm{w}$}, simplifying them to satisfy the orthogonally convex property required by our reward model for subsequent conversion into verification functions.
LLM-Guided Lemma Generation. The task of LLM Agent is to transform the existence conditions of the Steiner tree involving four specific points into a form that can be used for verification function. The process proceeds as follows:
@@ 4.3 Iterative Process
In this subsection, we describe the iterative closed-loop interaction between the LLM Agent and the reward model. While self-reflection mechanisms have proven pivotal in solving contest(Jaech et al., 2024; Singh et al., 2025; Lai et al., 2025)-level math problems—enabling models to review generation history, detect errors, and iteratively refine accuracy—they are not directly applicable here. In our system, the objective is not to solve the Gilbert-Pollak conjecture directly, but to synthesize intermediate lemmas. Consequently, generic reflection strategies are insufficient. Instead, we design a targeted supervisory signal that explicitly identifies areas for improvement, guiding the LLM to overcome specific bottlenecks during the refinement process.
@@
Verification of the Proof Correctness. We ensure the reliability of our results through a dual-verification strategy. First, the LLM acts solely as a reasoning engine, offloading all validation tasks to Mathematica to perform exact computation. This significantly reduces hallucinations in the generated lemmas. Second, we performed independent manual verification of the generated proofs to confirm that the overall logical argument is sound. Consequently, the final lower bound is a standard mathematical proof that stands independent of the generative AI pipeline.
@@ 5 Experiments and Results
Lemma Quality. Across approximately $100$ rounds of experiments, more than $80\%$ of LLM-proposed lemmas are effective when instantiated as verification functions. This high success rate highlights the advantages of operating within narrow reasoning spaces and using a localized, lemma-driven approach, rather than attempting end-to-end optimization over the full parameter space.
Robustness Across LLM Backbones. We conducted experiments using GPT-5 and Gemini 3 Pro. In all cases, the LLMs successfully proposed valid lemmas, and after roughly a dozen iterative rounds, the system consistently converged to a Steiner ratio lower bound of approximately $0.855$. These results demonstrate that the framework is robust and largely insensitive to the choice of LLM backbone.
@@
Figure 9: Illustration of the ratio change throughout the iteration over steps and compasion with SOTA result of previous works.
Typical Trajectory and Final Result. Figure 9 summarize a representative trajectory of the system. Through iterative interaction between the LLM agent and the reward model, the lower bound progressively improves and stabilizes at $0.8559$. Importantly, this bound has been rigorously verified across the entire parameter space, and all lemmas proposed along this trajectory have been manually checked for correctness, ensuring the validity of the overall proof independent of the specific generative AI pipeline. Across this trajectory, the LLM consumed approximately 0.50.5M tokens ($\sim 35$K per round) and $4.6$ hours of reasoning, while computations of the reward model required roughly $11.7$ hours. The detailed content of each step of improvement can be found in the Appendix C.

### v2

@@ 1 Introduction
Recently, Large Language Models (LLMs) have demonstrated remarkable results in solving contest-level mathematics (Achim et al., 2025; Chen et al., 2025; Chervonyi et al., 2025). A natural next step is to explore whether these capabilities can be extended to tackling unsolved mathematical conjectures, such as the Gilbert-Pollak Conjecture. However, naïvely prompting LLMs to address such tasks often yields incorrect proofs with subtle logical gaps and fundamental misunderstandings of key concepts. Unlike canonical definitions commonly found in textbooks, research-level concepts appear only sparsely in training data. Compared to well-structured competition problems, open research problems exhibit a vastly larger and less constrained search space. As a result, models struggle to identify valid proof strategies and become substantially more prone to hallucination.
@@
Reproducible codes and the final proof can be found in https://github.com/keyisi2006/Steiner-Ratio. If you are a mathematician interested solely in the mathematical results presented in this paper, you may proceed directly to Appendix H, where we provide a simplified version that excludes content related to LLMs.
@@ 2.2 Mathematical Framework for Proving Steiner Ratio Lower Bound
In this section, we briefly describe the mathematical framework for proving lower bound for Steiner ratio. Many of the mathematical details can be found in Appendix B.1.
We first give a very brief and informal description. The framework for establishing the Steiner ratio lower bound is mathematical induction (Du & Hwang, 1983). Assuming that for any set of points with cardinality at most n1n-1, the Steiner ratio is lower-bounded by $\rho$. To extend this to a set $V$ of $n$ points, the core strategy involves a reduction argument: identifying a specific subset $V^{\prime}\subset V$ of size nrn-r (where $r\geq 1$) such that the marginal gain in the Steiner tree length satisfies $\Delta L_{S}/\Delta L_{m}\geq\rho$. Here, $\Delta L_{S}=L_{S}(V)-L_{S}(V^{\prime})$ and $\Delta L_{m}=L_{m}(V)-L_{m}(V^{\prime})$ denote the reduction in the lengths of the Steiner Minimal Tree and the Minimum Spanning Tree, respectively. Since $|V^{\prime}|\leq n-1$, the inductive hypothesis guarantees $L_{S}(V^{\prime})/L_{m}(V^{\prime})\geq\rho$. Combining this with the marginal condition $\Delta L_{S}\geq\rho\Delta L_{m}$ completes the inductive step, proving $L_{S}(V)/L_{m}(V)\geq\rho$.
@@
We refer to the tuple $\tau=(V^{*},S^{-},S^{+},t^{*})$ arising from this pruning procedure as a splitting; a formal definition is provided in Appendix B.1.2.
Figure 4: Illustration of Pruning Process. We try to prune $V^{*}=\{B\}$. Black dashed lines are $S^{-}$. Black solid lines are $S^{+}$. Green line is $t^{*}$. Yellow points are Steiner points in $S^{+}$.
In this pruning process, the decrease in the Steiner tree length is given by $\Delta L_{S}\geq L_{S^{-}}-L_{S^{+}}$, while the increase in the spanning tree length is given by $\Delta L_{m}\leq L_{t^{*}}$. Here, $L_{X}$ denotes the total edge length of graph $X$. The requirement for the induction to hold is that $\Delta L_{S}/\Delta L_{m}\geq\rho$, which is equivalent to $\rho\cdot L_{t^{*}}+L_{S^{+}}-L_{S^{-}}\leq 0$. If this condition holds, then combining it with the inductive hypothesis $L_{S}(V^{\prime})/L_{m}(V^{\prime})\geq\rho$ allows us to derive $L_{S}(V)/L_{m}(V)\geq\rho$, thereby completing the induction step. Therefore, we can reformulate the verification of a candidate lower bound $\rho$ for the Steiner ratio as the minimax condition in Theorem 4. Let $\mathcal{W}=[0,+\infty)^{n}$ denote the parameter space of edge lengths; see Definition 12 in Appendix B.1.1 for its precise construction.
@@
For any splitting $\tau=(V^{*},S^{-},S^{+},t^{*})$ and geometric configuration ${\bm{w}}\in\mathcal{W}$, we define the splitting function $F_{\tau}({\bm{w}},\rho)$ as: $F_{\tau}({\bm{w}},\rho)=\rho\cdot L_{t^{*}}+L_{S^{+}}-L_{S^{-}}.\vskip-5.0pt$ Let $\mathcal{F}=\{F_{\tau}:\tau\text{ is a splitting}\}$ denote the collection of all splitting functions. The lower bound $\rho_{\text{Steiner}}\geq\rho$ is established if: $\textstyle\max_{{\bm{w}}\in\mathcal{W}}\min_{F\in\mathcal{F}}F({\bm{w}},\rho)\leq 0.$ (1) We say that a value $\rho$ is feasible if the condition equation 1 holds.
In the following sections, we introduce an LLM-based mathematical framework designed to address this minimax problem.
@@ 3 LLM Math System for the Conjecture
Finally, we develop a bottleneck region-based reflection mechanism. Leveraging the reward signal and the verification process, we introduce an algorithm that automatically identifies key bottleneck regions limiting further improvements of the lower bound and feeds this information back to the LLM in natural language, guiding subsequent generations to overcome these bottlenecks. Experiments show that this signal significantly improves the effectiveness of LLMs, thereby addressing Challenge (3).
@@ 3.1 Reward Model
The general objective of the reward in the LLM is to evaluate the generated output. In our context, the reward is to determine the tightest lower bound for the Steiner ratio from the “generated proof”. Theorem 4 establishes that the statement “the Steiner ratio is lower bounded by $\rho$” is equivalent to verifying that $\max_{{\bm{w}}\in\mathcal{W}}\min_{F\in\mathcal{F}}F({\bm{w}},\rho)\leq 0$. However, directly evaluating this continuous-domain minimax problem is computationally intractable.
@@
A function $f:\mathcal{W}\to\mathbb{R}\cup\{+\infty\}$ is called a verification function if it satisfies the following two conditions: 1. Shape Constraint: The restriction of $f$ to any line parallel to a coordinate axis is unimodal (first non-increasing and then non-decreasing). 2. Bounding Constraint: There exists a splitting function (defined in Theorem 4) $g\in\mathcal{F}$ such that $g({\bm{w}})\leq f({\bm{w}})$, $\forall{\bm{w}}\in\mathcal{W}$. We denote the set of all verification functions by $\mathcal{F}_{\text{ver}}$.
@@
Let $H\subset\mathcal{W}$ be a bounded axis-aligned hyperrectangle and let $f$ be any function in $\mathcal{F}_{\text{ver}}$. A key property is that the global maximum of $f$ over $H$ is attained at one of its vertices. Consequently, if $\textstyle\max_{{\bm{w}}\in\mathcal{V}(H)}f({\bm{w}})\leq 0,$ (2) then $f({\bm{w}})\leq 0$ for all ${\bm{w}}\in H$, which implies that the underlying splitting function $g({\bm{w}})\leq 0$. Note that $\mathcal{V}(H)$ is the set of vertices of $H$.
Leveraging these theoretical properties of the verification function, we construct a reward model to compute the lower bound of the Steiner ratio. We present the core algorithm below, deferring the full pseudocode and detailed explanations to Appendix F.1.
@@
(1) Region Partitioning. We recursively partition the infinite parameter space into hyperrectangle regions. (2)Vertex Verification. For bounded regions, we select a verification function $f$ and invoke the Vertex Maximum Property. By verifying $f({\bm{w}})\leq 0$ at all vertices, we mathematically guarantee that the inequality holds for every point $\bm{w}$} within the region. (3) Monotonicity Check for Unbounded Regions. For unbounded regions, we first verify that $f$ is non-increasing along those unbounded dimensions, ensuring the maximum value occurs at the finite lower boundary (a “slice” of the region). We then apply vertex verification to this bounded slice. Details of the monotonicity check are provided in Appendix F.2. (4) Refinement via Subdivision. Unverified regions are subdivided along one dimension into two halves, and recursively attempt verification.
@@ 3.2 LLM Agent
We demonstrate that finding a verification function reduces to identifying two types of simpler geometric lemmas: (1) the Trapped Regular Point Lemma and (2) the Valid 4-Point Steiner Tree Lemma. (This equivalence is illustrated in Figure 8 and formally proven in Theorem 20, Appendix C.2). Within our framework, the LLM generates candidate lemmas in the form of structured code snippets. This approach significantly simplifies the generation process and facilitates rigorous verification. In the subsequent sections, we detail how we leverage the LLM to discover these lemmas.
@@ 3.2.1 Finding Trapped Regular Point Lemma
In this section, we detail the methodology for discovering Trapped Regular Point Lemmas using LLMs. We define two distinct categories of rules: Type A rules encode necessary properties that any Steiner Minimal Tree (SMT) must satisfy, while Type B rules specify sufficient conditions that guarantee the existence of an implicit regular point. Given the vast combinatorial space of rule interactions—which renders manual exploration intractable (see Appendix G for details)—we employ an LLM integrated with local tools (e.g., Mathematica). By iteratively validating hypotheses, the LLM discovers effective rule combinations and synthesizes conditions that imply regularity. These conditions are subsequently simplified by the model to derive novel geometric lemmas.
@@
These predicates correspond to algebraic inequalities on $\bm{w}$} (e.g., $C^{B}_{0,2,1}\iff(a_{2}\leq 1)\land(\sqrt{3}/2\cdot(1+a_{2})\leq 1)$). Full derivations are in Appendix C.1.
@@
LLM-Guided Lemma Generation. This module operationalizes Theorem 9 through the synthesis of valid constraints via an LLM agent. Here, we delineate the process of guiding the LLM to identify trapped regular point lemmas. Comprehensive details are available in Appendix G.
@@ 3.2.2 Finding Valid 4-Point Steiner Tree Lemma
We adopt the Valid 4-Point Steiner Tree Theorem from Du et al. (1987) to determine the existence of the (AB)(AB)-$(CD)$ topology. The theorem’s details are provided in Appendix B.2 as Theorem 16. The LLM agent’s task is to translate the geometric conditions of the theorem into algebraic constraints on the parameter space $\bm{w}$}, simplifying them to satisfy the orthogonally convex property required by our reward model for subsequent conversion into verification functions.
LLM-Guided Lemma Generation. The task of the LLM agent is to transform the existence conditions of the Steiner tree involving four specific points into a form that can be used for verification functions. The process proceeds as follows:
@@ 3.3 Iterative Process
In this subsection, we describe the iterative closed-loop interaction between the LLM agent and the reward model. While self-reflection mechanisms have proven pivotal in solving contest-level math problems (Jaech et al., 2024; Singh et al., 2025; Lai et al., 2026)—enabling models to review generation history, detect errors, and iteratively refine accuracy—they are not directly applicable here. In our system, the objective is not to solve the Gilbert-Pollak conjecture directly, but to synthesize intermediate lemmas. Consequently, generic reflection strategies are insufficient. Instead, we design a targeted supervisory signal that explicitly identifies areas for improvement, guiding the LLM to overcome specific bottlenecks during the refinement process.
@@
Verification of the Proof Correctness. We ensure the reliability of our results through a dual-verification strategy. First, the LLM acts solely as a reasoning engine, offloading all validation tasks to Mathematica to perform exact computations. This significantly reduces hallucinations in the generated lemmas. Second, we performed independent manual verification of the generated proofs to confirm that the overall logical argument is sound. Consequently, the final lower bound is a standard mathematical proof that stands independent of the generative AI pipeline.
@@ 4 Experiments and Results
Lemma Quality. Across approximately $100$ rounds of experiments, more than $80\%$ of LLM-proposed lemmas are effective (both mathematically correct and helpful in improving the lower bound for the current step) when instantiated as verification functions. This high success rate highlights the advantages of operating within narrow reasoning spaces and using a localized, lemma-driven approach, rather than attempting end-to-end optimization over the full parameter space.
Robustness Across LLM Backbones. We conducted experiments using GPT-5 and Gemini 3 Pro. In all cases, the LLMs successfully proposed valid lemmas, and after roughly a dozen iterative rounds, the system consistently converged to a Steiner ratio lower bound of approximately $0.8559$. These results demonstrate that the framework is robust and largely insensitive to the choice of LLM backbone.
@@
Figure 9: Illustration of the ratio change throughout the iteration over steps and comparison with best results of previous works.
Typical Trajectory and Final Result. Figure 9 summarize a representative trajectory of the system. Through iterative interaction between the LLM agent and the reward model, the lower bound progressively improves and stabilizes at $0.8559$. Importantly, this bound has been rigorously verified across the entire parameter space, and all lemmas proposed along this trajectory have been manually checked for correctness, ensuring the validity of the overall proof independent of the specific generative AI pipeline. Across this trajectory, the LLM consumed approximately 0.50.5M tokens ($\sim 35$K per round) and $4.6$ hours of reasoning, while computations of the reward model required roughly $11.7$ hours. The detailed content of each step of improvement can be found in the Appendix E.